# SpecEval: Evaluating Code Comprehension in Large Language Models via Program Specifications

### v1

@@ 5.1.2. Comparison Between Models
Model Metric DUB IEF IDS NMR NMS Overall Model Metric DUB IEF IDS NMR NMS Overall
GPT-4 $J_{jud}$ 0.153 0.110 0.065 0.144 0.165 0.455 GPT-3.5 $J_{jud}$ 0.323 0.337 0.362 0.358 0.348 0.419
$J_{sel}$ 0.120 0.107 0.101 0.163 0.141 $J_{sel}$ 0.303 0.263 0.239 0.366 0.277
$J_{inf}$ 0.818 0.825 0.860 0.714 0.805 $J_{inf}$ 0.535 0.484 0.586 0.594 0.657
$v_{prec}$ 0.630 0.699 0.775 0.623 0.719 $v_{prec}$ 0.490 0.499 0.445 0.495 0.527
$v_{rec}$ 0.518 0.523 0.586 0.460 0.539 $v_{rec}$ 0.399 0.399 0.345 0.422 0.422
Avg. 0.448 0.453 0.477 0.421 0.474 Avg. 0.410 0.396 0.395 0.447 0.446
Llama $J_{jud}$ 0.457 0.496 0.511 0.534 0.493 0.609 CodeLlama $J_{jud}$ 0.275 0.248 0.250 0.340 0.235 0.409
$J_{sel}$ 0.845 0.798 0.840 0.837 0.854 $J_{sel}$ 0.413 0.292 0.308 0.458 0.373
$J_{inf}$ 0.962 0.828 0.667 0.848 0.852 $J_{inf}$ 0.612 0.535 0.500 0.759 0.583
$v_{prec}$ 0.536 0.527 0.533 0.538 0.542 $v_{prec}$ 0.500 0.447 0.488 0.455 0.502
$v_{rec}$ 0.346 0.364 0.349 0.338 0.337 $v_{rec}$ 0.326 0.339 0.325 0.314 0.345
Avg. 0.629 0.603 0.580 0.619 0.616 Avg. 0.425 0.372 0.374 0.465 0.408
Deepseek -Coder $J_{jud}$ 0.389 0.376 0.362 0.396 0.374 0.441 Magicoder $J_{jud}$ 0.526 0.563 0.521 0.521 0.523 0.485
$J_{sel}$ 0.236 0.284 0.210 0.252 0.230 $J_{sel}$ 0.597 0.524 0.580 0.579 0.529
$J_{inf}$ 0.643 0.709 0.596 0.679 0.720 $J_{inf}$ 0.632 0.511 0.425 0.571 0.570
$v_{prec}$ 0.517 0.506 0.488 0.492 0.501 $v_{prec}$ 0.434 0.418 0.452 0.463 0.450
$v_{rec}$ 0.399 0.426 0.413 0.402 0.422 $v_{rec}$ 0.319 0.348 0.345 0.355 0.371
Avg. 0.437 0.460 0.414 0.444 0.449 Avg. 0.502 0.473 0.465 0.498 0.489
5.2. RQ2: Counterfactual Analysis
@@
The Jaccard Distances and Average Variances of different models on different tasks are listed in Table 2. As described in Section 3.3, both denote the perturbations’ impact. A Higher Jaccard Distance or Average Variance indicates a higher impact imposed by the perturbation. Generally, all models exhibit performance variance to different degrees, indicating that the semantics LLMs learn are interfered with by the perturbations. It can be observed that the impact of perturbations varies differently among different tasks, and the impact tends to be stronger in the tasks where the models underperform. For instance, as described in Section 5.1, GPT-4 significantly outperforms all the other models in terms of Judgement and Selection task, where GPT-4 also possesses the lowest $J_{jud}$ and $J_{sel}$, showing performance consistency over all types of perturbations on these two tasks. Also, the performance of all models showcases a significant decline in the Infilling tasks, and correspondingly, the $J_{inf}$ metrics of all models are also significantly higher than that of other tasks. The models that perform relatively better in Infilling, such as Magicoder and GPT-3.5, also showcase stronger resistance against the perturbations with relatively lower $J_{inf}$.
We further measure the average impact of different perturbations on each model across 5 tasks. It is calculated by the average of all impact metrics of each perturbation across 5 tasks. The Avg. row presents the average impact of different perturbations in Table 2. We can observe that different models are sensitive to different types of perturbations. For instance, GPT-4 is sensitive to Independent Swap (i.e., IDS in Table 2) and Name shuffle (i.e., NMS), with an exceptionally high variance of 0.477 and 0.474 in these categories. GPT-3.5 is relatively sensitive to Name Random (i.e., NMR) and Name shuffle, Deepseek-Coder is relatively weak at If-else Flip (i.e., IEF), and Magicoder is relatively weak at Def-use Break (i.e., DUB). Concerning the average impact on all models, Name Random achieves the highest average impact of 0.482, followed by Name Shuffle, with an average impact of 0.480. These two types of perturbations showcase the most decisive influence on model outputs. The following perturbation is Def-use Break, with a slightly lower impact of 0.475. The other two types of perturbation, If-else Flip and Independent Swap, exhibit a relatively lower impact of 0.460 and 0.451, respectively.
We also calculated the overall impact of all perturbations on each model (presented in the columns starting with Overall in Table 2). Across all subject LLMs, CodeLlama showcases minimal performance variance against all types of perturbations, with the average impact on it being 0.409. The following model is GPT-3.5, with a slightly higher performance variance of 0.419. In general, the two leading models exhibit better resistance against perturbations. Other models demonstrate higher performance variance and more sensitivity to the perturbations.
Answer to RQ2: All LLMs exhibit sensitivities of different levels towards the perturbations, indicating that perturbations can interfere with the semantics learned. The impact of perturbations varies differently among different tasks, with a tendency to be stronger in the tasks where the models underperform.
Table 3. Progressive Consistency Score of each model on all perturbation categories.
Model Original Def-use Break If-else Flip Independent Swap Name Random Name Shuffle
GPT-3.5 0.276 0.272 0.263 0.266 0.240 0.237
GPT-4 0.262 0.260 0.229 0.248 0.238 0.243
Llama 0.124 0.121 0.127 0.120 0.135 0.124
CodeLlama 0.169 0.174 0.159 0.182 0.159 0.169
Deepseek-Coder 0.218 0.193 0.214 0.223 0.208 0.206
Magicoder 0.214 0.237 0.249 0.246 0.257 0.238
5.3. RQ3: Progressive Consistency Analysis
The Progressive Consistency Scores (PCS) of each model on all perturbation categories are presented in Table 3. Generally, all LLMs achieve a relatively low PCS under 0.3, indicating that all LLMs exhibit limited consistency between the tasks involved in the experiments. Considering the relatively low performance of all subject LLMs on the third task Infilling mentioned in Section 5.1, this phenomenon is reasonable. Underperformance on the third task indicates that LLMs can only solve at most two tasks in a row for the majority of programs. The model exhibiting the highest consistency is GPT-3.5, with an average PCS above 0.25 among 4 out of 6 categories. The following model is GPT-4, with a small step behind GPT-3.5 in terms of PCS. The PCS of the above two models is around 0.25, indicating that most of the programs succeed in handling two tasks consecutively. Deepseek-Coder and Magicoder form the second tier, both achieving an average PCS above 0.2 in most categories. Llama and CodeLlama achieve a relatively lower PCS below 0.2, exhibiting the least consistency among all models.
This observation shows that code LLMs do not necessarily exhibit better progressive consistency than general LLMs, despite their performance on each task mentioned in Section 5.1. Since GPT-4 and GPT-3.5 are closed-source commercial LLMs trained with larger model scale and possibly more immense data corpora, they exhibit better text comprehension and logical reasoning abilities, which can be leveraged in our tasks, especially the first task Judgement and the second task Selection. This leads to the solid and excelling performance of these two models on the former two tasks, guaranteeing a stable PCS above 0.25. Despite the impressive abilities showcased by open-source models such as Magicoder in the latter two tasks, Infilling and Generation, the tasks are relatively more challenging and cannot be easily handled in a row. Limited Progressive Consistency Scores can be earned through them, eventually leading to the gap behind GPT-4 and GPT-3.5.
In terms of consistency across different categories, we can witness declines within 3~4 out of 6 models for each perturbed category compared to Original category, but usually not significant. Nevertheless, in the perturbations with the strongest impact, i.e., Name Random and Name Shuffle, we can witness some decline in the consistency of GPT-3.5 and GPT-4. However, Magicoder achieves its highest PCS in Name Random across all categories, indicating that the impact of perturbations on the performance consistency varies across different models.
Answer to RQ3: The LLMs exhibit limited progressive consistency, indicating that they are not reasoning in a progressive manner (top-down or bottom-up) as humans do. GPT-4 and GPT-3.5 showcase a relatively better consistency due to their excelling performance in Judgement and Selection tasks. Future research remains focused on enhancing the performance of LLMs in tasks involving continuous and contextually linked reasoning.
@@ 6.1. Implication
The potential implications of this work cover a list of areas. Concerning the Evaluations on LLMs, semantic-preserving perturbations are proved to have a non-negligible impact on the performance of LLMs, which is worth attention when designing further benchmarks or evaluation frameworks. The resistance against perturbations should also be emphasized when developing new language models. Regarding Code Comprehension with LLMs, the assistance of program specifications in natural language or formal languages deserves to be noted. In terms of Code Completion and Generation with LLMs, the possibility can be explored concerning the utilization of specifications to assist the quality assurance of LLM-generated code. Concerning the Software Development process, the importance of all types of specifications (including documents and comments) should not be overlooked since they convey rich information about code and improve its readability and maintainability. As for Education, it is possible that program specifications in natural or formal languages can assist students in understanding example code during the teaching process.
@@ 6.2. Greedy Decoding for Model Generation
Following previous works (Liu et al., 2023; Du et al., 2024; Liu et al., 2024a), we adopted greedy decoding (top_k=1 and temperature=0) for all subject LLMs in the experiments. This is to guarantee the determinism in the sampling process of LLMs and the reproducibility of the final results. Nevertheless, some research within the same domain (Gu et al., 2024; Chen et al., 2024) adopted random sampling for evaluation, which is also a choice. Currently, research has emerged to compare the performance of LLMs under greedy decoding and other decoding strategies. Cobbe et al. (Cobbe et al., 2021) and Hendrycks et al. (Hendrycks et al., 2021) show that greedy decoding is preferred when math reasoning problems are to solve. Song et al. (Song et al., 2024) claim that greedy decoding is generally more effective for most tasks, especially reasoning tasks and coding problems. Since the specification-related tasks involved in this work are generally about reasoning and formulating formal statements, greedy decoding tends to be the more suitable choice.
6.3. Threats to Validity
@@
Internal Threats. First, experimental results may be affected by the prompts utilized to communicate with LLMs. We follow the prompt structure of Chen et al. (Chen et al., 2024) when designing the prompts. The effect of prompt strategies on the model performance will be studied in future works. Second, the base datasets, based on which we formulate the adapted benchmark, face potential data leakage when using the LLMs for evaluation. For programs in Frama-C-Problems (github, 2024) and SV-COMP (sosy lab, 2024), the ground-truth JML specifications involved are manually crafted by experts without risks of data leakage. As for SpecGenBench (Ma et al., 2024), the dataset was publicly released in March 2024, whereas the versions of models we adopted were all released prior to this point in time. Thus, using our selected LLMs for evaluation is without data leakage issues. However, if the selected models are after March 2024, there may be potential data leakage risks if the data is used for training.
@@
External Threats. The potential threat lies in the specification verifier. OpenJML (Cok, 2011) provides static verification and runtime checking for validation. Static verification requires sufficient supporting information (i.e., other strong specifications) and fails easily if not given enough, introducing serious bias. Therefore, runtime checking, with more stability and reliability, is adopted instead.
@@
In this paper, we present SpecEval, a novel black-box framework evaluating the code comprehension ability in Large Language Models via program specifications. Leveraging the feature that specifications cover abundant execution traces and rich semantic information, we adopt them as an effective representation of program semantics. Four specification-related tasks with sequential dependencies are designed to assess the performance of LLMs. Counterfactual Analysis and Progressive Consistency Analysis are adopted to analyze the performance consistency of LLMs over different dimensions. Experimental results show that LLMs showcase a limited ability to fully articulate program semantics with formal specifications. Counterfactual Analysis and Progressive Consistency Analysis also reveal the inconsistency in their performance over different dimensions.

### v2

@@ 5.2. RQ2: Effectiveness of the Few-shot Setting
Model Metric DUB IEF IDS NMR NMS Avg. Model Metric DUB IEF IDS NMR NMS Avg. Model Metric DUB IEF IDS NMR NMS Avg.
GPT-4 $J_{jud}$ 0.153 0.110 0.065 0.144 0.165 0.183 GPT-3.5 $J_{jud}$ 0.323 0.337 0.362 0.358 0.348 0.307 Llama $J_{jud}$ 0.457 0.496 0.511 0.534 0.493 0.573
$J_{sel}$ 0.120 0.107 0.101 0.163 0.141 $J_{sel}$ 0.303 0.263 0.239 0.366 0.277 $J_{sel}$ 0.845 0.798 0.840 0.827 0.854
$J_{inf}$ 0.114 0.058 0.057 0.165 0.153 $J_{inf}$ 0.188 0.130 0.163 0.298 0.325 $J_{inf}$ 0.875 0.694 0.548 0.806 0.857
$v_{prec}$ 0.186 0.126 0.179 0.152 0.174 $v_{prec}$ 0.240 0.238 0.205 0.227 0.226 $v_{prec}$ 0.318 0.289 0.300 0.288 0.300
$v_{rec}$ 0.362 0.348 0.399 0.421 0.404 $v_{rec}$ 0.463 0.462 0.407 0.483 0.443 $v_{rec}$ 0.513 0.431 0.461 0.482 0.510
Avg. 0.187 0.150 0.160 0.209 0.207 Avg. 0.303 0.286 0.275 0.346 0.324 Avg. 0.602 0.542 0.532 0.587 0.603
CodeLlama $J_{jud}$ 0.275 0.262 0.271 0.324 0.213 0.391 Deepseek Coder $J_{jud}$ 0.389 0.376 0.362 0.396 0.374 0.340 Magicoder $J_{jud}$ 0.526 0.563 0.521 0.521 0.523 0.496
$J_{sel}$ 0.373 0.222 0.303 0.459 0.329 $J_{sel}$ 0.236 0.284 0.210 0.252 0.230 $J_{sel}$ 0.597 0.524 0.580 0.579 0.529
$J_{inf}$ 0.544 0.451 0.455 0.617 0.533 $J_{inf}$ 0.318 0.254 0.203 0.375 0.338 $J_{inf}$ 0.621 0.494 0.422 0.583 0.616
$v_{prec}$ 0.408 0.349 0.364 0.357 0.379 $v_{prec}$ 0.318 0.289 0.300 0.288 0.300 $v_{prec}$ 0.376 0.325 0.337 0.335 0.323
$v_{rec}$ 0.416 0.470 0.407 0.465 0.521 $v_{rec}$ 0.513 0.431 0.461 0.482 0.510 $v_{rec}$ 0.477 0.491 0.482 0.513 0.548
Avg. 0.403 0.351 0.360 0.444 0.395 Avg. 0.355 0.327 0.307 0.358 0.351 Avg. 0.519 0.480 0.468 0.506 0.508
5.3. RQ3: Counterfactual Analysis
@@
The Jaccard Distances and Average Variances of different models on different tasks are listed in Table 3, both reflecting the perturbations’ impact as described in Section 3.2. A higher value of these metrics indicates a higher impact imposed by the perturbation. Generally, all models exhibit performance variance to different degrees, indicating that the semantics LLMs learn are interfered with by the perturbations. Across all models, GPT-4 showcases minimal performance variance against all types of perturbations, with an average impact of 0.183, followed by GPT-3.5, with a slightly higher performance variance of 0.307. Among the open-source models, Deepseek-Coder exhibits the strongest consistency, with the average variance being 0.340. Other models demonstrate a higher performance variance and more sensitivity to perturbations. Also, the impact of perturbations varies among different tasks, tending to be stronger in the tasks where the models underperform. For instance, as described in Section 5.1, GPT-4 significantly outperforms all the other models in most circumstances, where it also possesses the lowest performance variance, showing the best consistency over all types of perturbations. Also, all models yield relatively lower Recall in the Generation task, and correspondingly, the $v_{rec}$ metrics of most models are also significantly higher than others.
Figure 4. The infilling task for program IntSquare (both the Original and the Name Random category), along with the answers provided by GPT-3.5 and CodeLlama.
New Findings: Concerning the average impact of the commonly used perturbations, Name Random demonstrated the most significant impact, with 4 out of the 6 models showing the most sensitivity to it, followed by Def-use Break and Name shuffle, which also pose a notable influence on the models’ performance. The commonality among these three types of perturbations lies in their alteration of variable names, which disrupts the semantic information contained. To further illustrate this, two versions (Original and Name Random perturbed) of program IntSquare along with their Infilling tasks are provided in Figure 4. The models should reason about the constraints on a series of loop variables to complete the loop invariant on Line 11. For the Original version, both GPT-3.5 and CodeLlama completed the task successfully. However, when it comes to the Name Random perturbed version, CodeLlama failed to grasp the functionality of each variable effectively faced with the perturbed identifiers. It only replicated the loop invariant from Line 7, which coincidentally shares the same prefix with the target one, resulting in an incorrect answer. In contrast, GPT-3.5 was able to maintain the correct reasoning. As suggested by previous works (Miceli-Barone et al., 2023; Wang et al., 2023; Liu et al., 2023b), variable naming can affect model performance. Intuitively, identifiers can carry semantic information such as variable functionalities, helping models in semantic comprehension. Irregular variable names reduce the semantic information that models can learn from them, causing fluctuations in model performance. Furthermore, this case reveals the tendency of LLMs to imitate the prompts given blindly when being unable to reason effectively, despite the potential inapplicability of these contents. Compared to these perturbations, If-else Flip and Independent Swap show a relatively lower impact since they have less interference with the approaches to acquiring program semantics. This is a novel finding compared to prior works (Gu et al., 2024; Chen et al., 2024a), since they did not conduct fine-grained Counterfactual Analysis incorporating semantic-preserving perturbations for LLMs, through which we discover the implicit semantic information embedded in identifiers and the impact of perturbing them on model performance.
Answer to RQ3: Perturbations can interfere with the semantics learned. Models tend to exhibit stronger sensitivity to perturbations in the tasks where they underperform. Among the semantic-preserving perturbations, the most influential ones are those that interfere with the identifiers. Since identifiers can carry semantic information, perturbations disrupting them can pose a greater impact on the models.
@@ 6.1. Implication
The implications of the intuitions described in Section 5 cover multiple parties involved with LLMs. For LLM developers, more refined datasets and fine-tuning techniques should be prepared to improve LLMs’ resistance against perturbations and extraneous information. For LLM testers, it is necessary to incorporate adversarial test cases with misleading information to thoroughly evaluate model robustness. For LLM-based tool developers, applying validation and error-correction mechanisms on model output is crucial to improving tool reliability. For software developers utilizing LLMs, the model prompts should be cautiously constructed to avoid interference from extraneous information that may disrupt the models’ reasoning. Except for those directly associated areas, in terms of code completion and generation with LLMs, specifications can assist the quality assurance of LLM-generated code. Concerning the software development process, all types of specifications (including documents and comments) deserve attention since they improve code readability and maintainability.
@@ 6.2. Greedy Decoding for Model Generation
Following previous works (Liu et al., 2023a; Du et al., 2024; Liu et al., 2024a), we adopted greedy decoding (top_k=1 and temperature=0) for all LLMs in the experiment to guarantee the determinism in the sampling process of LLMs and the reproducibility of the final results. Nevertheless, some research within the same domain (Gu et al., 2024; Chen et al., 2024a) adopted random sampling for evaluation, which is also a choice. Currently, research has emerged to compare the performance of LLMs under greedy decoding and other decoding strategies. Cobbe et al. (Cobbe et al., 2021) and Hendrycks et al. (Hendrycks et al., 2021) show that greedy decoding is preferred when math reasoning problems are to solve. Song et al. (Song et al., 2024) claim that greedy decoding is generally more effective for most tasks, especially reasoning tasks and coding problems. Since the specification-related tasks involved in this work are generally about reasoning and formulating formal statements, greedy decoding tends to be the more suitable choice.
6.3. Threats to Validity and Limitations
@@
Internal Threats. First, experimental results may be affected by the prompts utilized to communicate with LLMs. We follow the prompt structure of prior work. (Chen et al., 2024a). The effect of prompt techniques on the model performance will be studied in future works. Second, the base datasets, based on which we formulate the adapted benchmark, face potential data leakage when using the LLMs for evaluation. For programs in Frama-C-Problems (github, 2024) and SV-COMP (sosy lab, 2024), the ground-truth JML specifications involved are manually crafted by experts without risks of data leakage. As for SpecGenBench (Ma et al., 2024), the dataset was publicly released in March 2024, whereas the versions of models we adopted were all released before this point in time. Thus, using our selected LLMs for evaluation is without data leakage issues. The last threat lies in the specification-related knowledge required by the experimental tasks, which some models may lack, potentially leading to under-performance in the tasks. To tackle this, we adopt the few-shot learning technique so that the models can learn the fundamental syntax and semantics of JML from the examples given, which is effective as is analyzed in Section 5.2.
@@
External Threats. The external threat lies in the specification verifier. OpenJML (Cok, 2011) provides static verification and runtime checking for validation. Static verification requires sufficient supporting information (i.e., other strong specifications) and fails easily if not given enough, introducing serious bias. Therefore, runtime checking, with more stability and reliability, is adopted instead.
Limitations. The main limitation lies in the limited size of our dataset due to the difficulties in manually crafting reliable specifications. To compensate for this, we design multiple tasks while performing various equivalent perturbations on the existing programs, thus increasing the volume of experiments, improving the diversity of experimental content, and enhancing the reliability of the experimental results. Our benchmark involves a total of 4896 tasks for each subject model evaluated, whereas the number is 3152 for REval. We plan to employ existing specification generation techniques (Ma et al., 2024; Wen et al., 2024) to enlarge the benchmark in the future.
7. Related work
LLM-based Specification Generation. Recent works have been focusing on generating program specifications with LLMs. Janssen et al. (Janßen et al., 2024) investigated the ability of ChatGPT (OpenAI, 2024) to infer loop invariants for C programs. Pei et al. (Pei et al., 2023) utilized fine-tuning to enhance the performance of LLMs on specification generation tasks. Ma et al. (Ma et al., 2024) combined LLM conversations with mutations to generate specifications for Java programs. Wei et al. (Wen et al., 2024) adopted the program decomposition technique to assist the LLM-based specification generation process. For Rust code, Yang et al. (Yang et al., 2024a) utilized LLM agent networks to generate proofs automatically, whereas Chen et al. (Chen et al., 2024b) established a self-evolving cycle incorporating data synthesis and fine-tuning to enhance LLMs’ generation ability. Compared to these works, we aim to evaluate LLMs’ code comprehension abilities through specification-related tasks.
LLM Benchmarks. Some popular code generation benchmarks such as HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021) and SWE-bench (Jimenez et al., 2023) have been proposed to evaluate LLMs’ coding capabilities. Other benchmarks targeting code comprehension with different downstream tasks have also been established. CRUXEval (Gu et al., 2024) is a dataset focusing on input/output prediction tasks for simple Python functions to evaluate the code reasoning and execution abilities of LLMs. Inspired by CRUXEval, REval (Chen et al., 2024a) further assesses LLMS’ code reasoning ability through four different tasks concerning dynamic execution prediction. Compared to CRUXEval and REval, our framework utilizes program specifications as a carrier of overall program semantics to evaluate LLMs’ general understanding on programs’ essential behaviors and functionalities, rather than reasoning about program states within a single execution trace. This novel perspective leads us to new findings such as the deficiency of LLMs in synthesizing the semantics of multiple control flow structures as described in Section 5.1. Additionally, some other works incorporating formal method tasks to evaluate LLM ability have been proposed. He et al. (He et al., 2024) introduce an LLM maturity model by assessing LLMs’ performance on post-condition generation tasks. Compared to this work, we focus on a wider variety of specification-related tasks and a broader spectrum of specification types (including pre-conditions, post-conditions and loop invariants) to conduct a more comprehensive evaluation. Cao et al. (Cao et al., 2025) constructed an extensive dataset to investigate model performance improvements on formal methods tasks brought about by fine-tuning. Compared to this work, we focus on the program semantics learned by LLMs by assessing the formal specifications generated, while incorporating Counterfactual Analysis to study the performance variance of LLMs when faced with semantic-preserving perturbations, making further discoveries regarding the semantic information embedded within identifiers as described in Section 5.3.
@@ 7. Related work
We present SpecEval, a novel black-box framework evaluating the code comprehension ability in Large Language Models via program specifications. Leveraging the feature that specifications cover abundant execution traces and rich semantic information, we adopt them as an effective representation of program semantics. Four specification-related tasks are designed to assess the performance of LLMs. Counterfactual Analysis is further designed to assess the sensitivity of LLMs to semantic-preserving perturbations. Experimental results show that LLMs showcase a limited ability to fully articulate program semantics with formal specifications, with performance inconsistency when confronted with perturbations.